I need to find the x and y intercepts for the equation.
4x-3y=9
Then I have to graph it.

Answer by mananth(16946)

4x-3y=9
3y=4x-9
/3
y=(4/3)x-3
when x= 0,y=-3 ( 0 , -3 )
when y =0,x=2.25( 9/4 , 0 )

m.ananth@hotmail.ca